a total of 27000 was invested, part of it at 10% and part at 12%. the total yield was 2990. how much was invested at each rate?
Answer by rfer(16322)

.1x+.12(27000-x)=2990
.1x+3240-.12x=2990
-.02x=-250
x=$12500
27000-x=$14500